# From Age-Structured Trophic Networks to Applied Control : Stabilization and Harvesting Strategies for Non-Transitive Competition and the Dynamics of Mosquitoes

### v1

@@ 3.2.2. Stability for $\psi_{i}\neq 0$
(3.137) $\underbrace{\begin{cases}\dot{\eta}_{1}=u^{*}-u-\phi_{2}(\eta_{2}),\\ \dot{\eta}_{2}=-\phi_{3}(\eta_{3}),\\ \dot{\eta}_{3}=-\phi_{1}(\eta_{1})\end{cases}}_{\psi_{i}\equiv 0},\;\underbrace{\begin{cases}\dot{\eta}_{1}=u^{*}-u-\hat{\phi}_{2},\\ \dot{\eta}_{2}=-\hat{\phi}_{3},\\ \dot{\eta}_{3}=-\hat{\phi}_{1}\end{cases}}_{\psi_{i}\neq 0}\;\text{with the following state variables :}\;\begin{cases}\eta_{2},\\ z_{1}=\eta_{3}-\alpha_{1}(\eta_{2}),\\ z_{2}=\eta_{1}-\alpha_{2}(z_{1}),\end{cases}$
(3.138) $\begin{aligned} u&=u^{*}+c_{3o}\phi_{1}(z_{2})-\theta\frac{c_{2}}{\lambda_{1}}\phi^{2}_{3}(z_{1})+(c_{2}-1-\theta)\phi_{3}(z_{1})+\frac{c_{2}}{\lambda_{1}}\phi_{3}(z_{1})\phi_{1}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-(c_{1}+1)\phi_{2}(\eta_{2})\\ \ &+\frac{\theta\phi_{3}(z_{1})}{\phi_{1}(z_{2})}\left(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+\,\phi_{3}(z_{1})\right).\end{aligned}$
@@
(3.139) $\dot{V}_{3}(\eta_{2},z_{1},z_{2})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z_{1})-c_{3}\phi^{2}_{1}(z_{2}),\quad\psi_{i}\equiv 0.$
(3.140) $\dot{V}_{3}(\eta_{2},z_{1},z_{2})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z_{1})-c_{3}\,\phi_{1}^{2}(z_{2})+A_{1}\left(\hat{\phi}_{1}-\phi_{1}(\eta_{1})\right)+A_{2}\,\left(\hat{\phi}_{2}-\phi_{2}(\eta_{2})\right)+A_{3}\left(\hat{\phi}_{3}-\phi_{3}(\eta_{3})\right),\quad\psi_{i}\neq 0.$
@@
(3.141) $\underbrace{\begin{cases}\dot{\eta}_{1}=u^{*}-u-\phi_{2}(\eta_{2}),\\ \dot{\eta}_{2}=-\phi_{3}(\eta_{3}),\\ \dot{\eta}_{3}=-\phi_{4}(\eta_{4}),\\ \dot{\eta}_{4}=-\phi_{1}(\eta_{1}),\end{cases}}_{\psi_{i}\equiv 0},\;\underbrace{\begin{cases}\dot{\eta}_{1}=u^{*}-u-\hat{\phi}_{2},\\ \dot{\eta}_{2}=-\hat{\phi}_{3},\\ \dot{\eta}_{3}=-\hat{\phi}_{4},\\ \dot{\eta}_{4}=-\hat{\phi}_{1},\end{cases}}_{\psi_{i}\neq 0}\;\text{with the following state variables:}\;\begin{cases}\eta_{2},\\ z_{1}=\eta_{3}-\alpha_{1}(\eta_{2}),\\ z_{2}=\eta_{4}-\alpha_{2}(z_{1}),\\ z_{3}=\eta_{1}-\alpha_{3}(z_{2}),\end{cases}$
(3.142) $\begin{aligned} u&=u^{*}+c_{40}\phi_{1}(z_{3})-\theta\frac{c_{3}}{\lambda_{1}}\phi^{2}_{4}(z_{2})+(c_{3}-1-\theta)\phi_{4}(z_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+\frac{c_{3}}{\lambda_{1}}\phi_{4}(z_{2})\phi_{1}(z_{3})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})-(c_{2}-1)\phi_{3}(z_{1})\\ &+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z_{1})+\frac{\theta}{\phi_{1}(z_{3})}\left(\phi^{2}_{4}(z_{2})-c_{1}\phi_{2}(\eta_{2})\phi_{4}(z_{2})+\phi^{2}_{3}(z_{1})\right)\\ &\frac{\theta\phi_{3}(z_{1})}{\phi_{1}(z_{3})}\left(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+(c_{2}-2)\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi^{2}_{4}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{4}(z_{2})\right),\end{aligned}$
@@
(3.143) $\dot{V}_{4}(\eta_{2},z_{1},z_{2},z_{3})=-\theta c_{1}\phi^{2}_{2}(\eta_{2})-\theta c_{2}\phi^{2}_{3}(z_{1})-\theta c_{3}\phi^{2}_{4}(z_{2})-c_{4}\phi^{2}_{1}(z_{3}),\quad\psi_{i}\equiv 0.$
(3.144) $\begin{aligned} \dot{V}_{4}(\eta_{2},z_{1},z_{2},z_{3})&=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z_{1})-\theta c_{3}\,\phi_{4}^{2}(z_{2})-c_{4}\,\phi_{1}^{2}(z_{3})+A_{1}\left(\hat{\phi}_{1}-\phi_{1}(\eta_{1})\right)+A_{2}\,\left(\hat{\phi}_{2}-\phi_{2}(\eta_{2})\right)+A_{3}\left(\hat{\phi}_{3}-\phi_{3}(\eta_{3})\right)\\ &+A_{4}\left(\hat{\phi}_{4}-\phi_{4}(\eta_{4})\right),\quad\psi_{i}\neq 0.\end{aligned}$
@@ 3.3. Toward generalization
We now extend the control design and Lyapunov based method to the general case $N\geq 3$ species, specifying the recursive form of the control laws and Lyapunov functions adapted to the cyclic coupling. Let us consider the following system, which describes the dynamics of non-transitive competition among $N+1$ species as described by the following system
@@
(3.146) $\underbrace{\begin{cases}\dot{\eta}_{1}=u^{*}-u-\phi_{2}(\eta_{2}),\\ \dot{\eta}_{2}=-\phi_{3}(\eta_{3}),\\ \dot{\eta}_{3}=-\phi_{4}(\eta_{4}),\\ \dot{\eta}_{4}=-\phi_{5}(\eta_{5}),\\ \dot{\eta}_{5}=-\phi_{6}(\eta_{6}),\\ \vdots\\ \vdots\\ \dot{\eta}_{N}=-\phi_{N+1}(\eta_{N+1}),\\ \dot{\eta}_{N+1}=-\phi_{1}(\eta_{1}),\end{cases}}_{\psi_{i}\equiv 0}\,\underbrace{\begin{cases}\dot{\eta}_{1}=u^{*}-u-\hat{\phi}_{2},\\ \dot{\eta}_{2}=-\hat{\phi}_{3},\\ \dot{\eta}_{3}=-\hat{\phi}_{4},\\ \dot{\eta}_{4}=-\hat{\phi}_{5},\\ \dot{\eta}_{5}=-\hat{\phi}_{6},\\ \vdots\\ \vdots\\ \dot{\eta}_{N}=-\hat{\phi}_{N+1},\\ \dot{\eta}_{N+1}=-\hat{\phi}_{1},\end{cases}}_{\psi_{i}\neq 0}\;\text{with the following state variables:}\;\begin{cases}\eta_{2},\\ z_{1}=\eta_{3}-\alpha_{1}(\eta_{2}),\\ z_{2}=\eta_{4}-\alpha_{2}(z_{1}),\\ z_{3}=\eta_{5}-\alpha_{3}(z_{2}),\\ z_{4}=\eta_{6}-\alpha_{4}(z_{3}),\\ \vdots\\ \vdots\\ z_{N-1}=\eta_{N+1}-\alpha_{N-1}(z_{N-2}),\\ z_{N}=\eta_{1}-\alpha_{N}(z_{N-1}).\end{cases}$
General formulation of the Lyapunov function and its derivative ($N\mapsto N+1$).
@@
(3.147) $\begin{aligned} V_{N}(\eta_{2},z_{1},\cdots,z_{N-1})&=\theta\lambda_{2}(e^{\eta_{2}}-1-\eta_{2})+\theta\sum_{i=1}^{N-2}\lambda_{i+2}(e^{z_{i}}-1-z_{i})+\lambda_{1}(e^{z_{N-1}}-1-z_{N-1}).\end{aligned}$
@@
(3.148) $\begin{aligned} \dot{V}_{N}(\eta_{2},z_{1},\dots,z_{N-1})={}&-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta\sum_{i=2}^{N-1}\,c_{\,i}\,\phi_{i+1}^{2}\bigl{(}z_{i-1}\bigr{)}-c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}\end{aligned},\quad\psi_{i}\equiv 0,$
@@
(3.149) $\begin{aligned} \dot{V}_{N}(\eta_{2},z_{1},\dots,z_{N-1})={}&-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta\sum_{i=2}^{N-1}\,c_{\,i}\,\phi_{i+1}^{2}\bigl{(}z_{i-1}\bigr{)}-c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}+\sum_{i=1}^{N}A_{i}\,\left(\hat{\phi}_{i}-\phi_{i}(\eta_{i})\right),\end{aligned}\quad\psi_{i}\neq 0,$

### v2

@@ 3.2.2. Stability for $\psi_{i}\neq 0$
(3.114) $\underbrace{\begin{cases}\dot{\eta}_{1}=u^{*}-u-\phi_{2}(\eta_{2}),\\ \dot{\eta}_{2}=-\phi_{3}(\eta_{3}),\\ \dot{\eta}_{3}=-\phi_{1}(\eta_{1})\end{cases}}_{\psi_{i}\equiv 0},\;\underbrace{\begin{cases}\dot{\eta}_{1}=u^{*}-u-\hat{\phi}_{2},\\ \dot{\eta}_{2}=-\hat{\phi}_{3},\\ \dot{\eta}_{3}=-\hat{\phi}_{1}\end{cases}}_{\psi_{i}\neq 0}\;\text{with the following state variables :}\;\begin{cases}\eta_{2},\\ z^{3}_{1}=\eta_{3}-\alpha_{1}(\eta_{2}),\\ z^{3}_{2}=\eta_{1}-\alpha_{2}(z^{3}_{1}),\end{cases}$
(3.115) $\begin{aligned} u&=u^{*}+c_{3o}\phi_{1}(z^{3}_{2})-\theta\frac{c_{2}}{\lambda_{1}}\phi^{2}_{3}(z^{3}_{1})+(c_{2}-1-\theta)\phi_{3}(z^{3}_{1})+\frac{c_{2}}{\lambda_{1}}\phi_{3}(z^{3}_{1})\phi_{1}(z^{3}_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})-(c_{1}+1)\phi_{2}(\eta_{2})\\ \ &+\frac{\theta\phi_{3}(z^{3}_{1})}{\phi_{1}(z^{3}_{2})}\left(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+\,\phi_{3}(z^{3}_{1})\right).\end{aligned}$
@@
(3.116) $\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z^{3}_{1})-c_{3}\phi^{2}_{1}(z^{3}_{2}),\quad\psi_{i}\equiv 0.$
(3.117) $\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z^{3}_{1})-c_{3}\,\phi_{1}^{2}(z^{3}_{2})+A_{1}\left(\hat{\phi}_{1}-\phi_{1}(\eta_{1})\right)+A_{2}\,\left(\hat{\phi}_{2}-\phi_{2}(\eta_{2})\right)+A_{3}\left(\hat{\phi}_{3}-\phi_{3}(\eta_{3})\right),\quad\psi_{i}\neq 0.$
@@
(3.118) $\underbrace{\begin{cases}\dot{\eta}_{1}=u^{*}-u-\phi_{2}(\eta_{2}),\\ \dot{\eta}_{2}=-\phi_{3}(\eta_{3}),\\ \dot{\eta}_{3}=-\phi_{4}(\eta_{4}),\\ \dot{\eta}_{4}=-\phi_{1}(\eta_{1}),\end{cases}}_{\psi_{i}\equiv 0},\;\underbrace{\begin{cases}\dot{\eta}_{1}=u^{*}-u-\hat{\phi}_{2},\\ \dot{\eta}_{2}=-\hat{\phi}_{3},\\ \dot{\eta}_{3}=-\hat{\phi}_{4},\\ \dot{\eta}_{4}=-\hat{\phi}_{1},\end{cases}}_{\psi_{i}\neq 0}\;\text{with the following state variables:}\;\begin{cases}\eta_{2},\\ z^{4}_{1}=\eta_{3}-\alpha_{1}(\eta_{2}),\\ z^{4}_{2}=\eta_{4}-\alpha_{2}(z^{4}_{1}),\\ z^{4}_{3}=\eta_{1}-\alpha_{3}(z^{4}_{2}),\end{cases}$
(3.119) $\begin{aligned} u&=u^{*}+c_{40}\phi_{1}(z^{4}_{3})-\theta\frac{c_{3}}{\lambda_{1}}\phi^{2}_{4}(z^{4}_{2})+(c_{3}-1-\theta)\phi_{4}(z^{4}_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+\frac{c_{3}}{\lambda_{1}}\phi_{4}(z^{4}_{2})\phi_{1}(z^{4}_{3})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z^{4}_{1})\phi_{4}(z^{4}_{2})-(c_{2}-1)\phi_{3}(z^{4}_{1})\\ &+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{3}(z^{4}_{1})+\frac{\theta}{\phi_{1}(z^{4}_{3})}\left(\phi^{2}_{4}(z^{4}_{2})-c_{1}\phi_{2}(\eta_{2})\phi_{4}(z^{4}_{2})+\phi^{2}_{3}(z^{4}_{1})\right)\\ &\frac{\theta\phi_{3}(z^{4}_{1})}{\phi_{1}(z_{3})}\left(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{4}_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})-\frac{c_{2}}{\lambda_{4}}\phi_{3}(z^{4}_{1})\phi_{4}(z^{4}_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+(c_{2}-2)\phi_{4}(z^{4}_{2})+\frac{c_{2}}{\lambda_{4}}\phi^{2}_{4}(z^{4}_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\phi_{4}(z^{4}_{2})\right),\end{aligned}$
@@
(3.120) $\dot{V}_{4}(\eta_{2},z^{4}_{1},z^{4}_{2},z^{4}_{3})=-\theta c_{1}\phi^{2}_{2}(\eta_{2})-\theta c_{2}\phi^{2}_{3}(z^{4}_{1})-\theta c_{3}\phi^{2}_{4}(z^{4}_{2})-c_{4}\phi^{2}_{1}(z^{4}_{3}),\quad\psi_{i}\equiv 0.$
(3.121) $\begin{aligned} \dot{V}_{4}(\eta_{2},z^{4}_{1},z^{4}_{2},z^{4}_{3})&=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z^{4}_{1})-\theta c_{3}\,\phi_{4}^{2}(z^{4}_{2})-c_{4}\,\phi_{1}^{2}(z^{4}_{3})+A_{1}\left(\hat{\phi}_{1}-\phi_{1}(\eta_{1})\right)+A_{2}\,\left(\hat{\phi}_{2}-\phi_{2}(\eta_{2})\right)+A_{3}\left(\hat{\phi}_{3}-\phi_{3}(\eta_{3})\right)\\ &+A_{4}\left(\hat{\phi}_{4}-\phi_{4}(\eta_{4})\right),\quad\psi_{i}\neq 0.\end{aligned}$
@@ 3.3. Toward generalization
We now extend the control design and Lyapunov based method to the general case $N\geq 3$ species, specifying the recursive form of the control laws and Lyapunov functions adapted to the cyclic coupling. Let us consider the following system, which describes the dynamics of non-transitive competition among $N$ species as described by the following system
@@
(3.123) $\underbrace{\begin{cases}\dot{\eta}_{1}=u^{*}-u-\phi_{2}(\eta_{2}),\\ \dot{\eta}_{2}=-\phi_{3}(\eta_{3}),\\ \dot{\eta}_{3}=-\phi_{4}(\eta_{4}),\\ \dot{\eta}_{4}=-\phi_{5}(\eta_{5}),\\ \dot{\eta}_{5}=-\phi_{6}(\eta_{6}),\\ \vdots\\ \vdots\\ \dot{\eta}_{N-1}=-\phi_{N}(\eta_{N}),\\ \dot{\eta}_{N}=-\phi_{1}(\eta_{1}),\end{cases}}_{\psi_{i}\equiv 0}\,\underbrace{\begin{cases}\dot{\eta}_{1}=u^{*}-u-\hat{\phi}_{2},\\ \dot{\eta}_{2}=-\hat{\phi}_{3},\\ \dot{\eta}_{3}=-\hat{\phi}_{4},\\ \dot{\eta}_{4}=-\hat{\phi}_{5},\\ \dot{\eta}_{5}=-\hat{\phi}_{6},\\ \vdots\\ \vdots\\ \dot{\eta}_{N-1}=-\hat{\phi}_{N},\\ \dot{\eta}_{N}=-\hat{\phi}_{1},\end{cases}}_{\psi_{i}\neq 0}\;\text{with the following state variables:}\;\begin{cases}\eta_{2},\\ z^{N}_{1}=\eta_{3}-\alpha_{1}(\eta_{2}),\\ z^{N}_{2}=\eta_{4}-\alpha_{2}(z^{N}_{1}),\\ z^{N}_{3}=\eta_{5}-\alpha_{3}(z^{N}_{2}),\\ z^{N}_{4}=\eta_{6}-\alpha_{4}(z^{N}_{3}),\\ \vdots\\ \vdots\\ z^{N}_{N-2}=\eta_{N}-\alpha_{N-2}(z^{N}_{N-3}),\\ z^{N}_{N-1}=\eta_{1}-\alpha_{N-1}(z^{N}_{N-2}).\end{cases}$
General formulation of the Lyapunov function and its derivative.
@@
(3.124) $\begin{aligned} V_{N}(\eta_{2},z^{N}_{1},\cdots,z^{N}_{N-1})&=\theta\lambda_{2}(e^{\eta_{2}}-1-\eta_{2})+\theta\sum_{i=1}^{N-2}\lambda_{i+2}(e^{z^{N}_{i}}-1-z^{N}_{i})+\lambda_{1}(e^{z^{N}_{N-1}}-1-z^{N}_{N-1}).\end{aligned}$
@@
(3.125) $\begin{aligned} \dot{V}_{N}(\eta_{2},z^{N}_{1},\dots,z^{N}_{N-1})={}&-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta\sum_{i=2}^{N-1}\,c_{\,i}\,\phi_{i+1}^{2}\bigl(z^{N}_{i-1}\bigr)-c_{N}\,\phi_{1}^{2}\bigl(z^{N}_{N-1}\bigr)\end{aligned},\quad\psi_{i}\equiv 0,$
@@
(3.126) $\begin{aligned} \dot{V}_{N}(\eta_{2},z^{N}_{1},\dots,z^{N}_{N-1})={}&-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta\sum_{i=2}^{N-1}\,c_{\,i}\,\phi_{i+1}^{2}\bigl(z^{N}_{i-1}\bigr)-c_{N}\,\phi_{1}^{2}\bigl(z^{N}_{N-1}\bigr)+\sum_{i=1}^{N}A_{i}\,\left(\hat{\phi}_{i}-\phi_{i}(\eta_{i})\right),\end{aligned}\quad\psi_{i}\neq 0,$
@@
(3.127) $\begin{aligned} u\;=&\;u^{*}+\frac{1}{\phi_{1}(z^{N}_{N-1})}\Bigg(c_{N}\,\phi_{1}^{2}(z^{N}_{N-1})+\theta\,c_{1}\,\phi_{2}^{2}(\eta_{2})+\theta\sum_{i=1}^{N-2}c_{i+1}\,\phi_{i+2}^{2}(z^{N}_{i})\Bigg)\\ &+\frac{\phi_{1}(z^{N}_{N-1})-\theta\,\phi_{N}(z^{N}_{N-2})}{\phi_{1}(z^{N}_{N-1})}\,\phi_{1}(\eta_{1})-\phi_{2}(\eta_{2})\\ &+\sum_{k=3}^{N}\frac{\displaystyle\theta\sum_{i=\max(1,k-3)}^{N-2}[(-1)^{\,i+\varepsilon_{k}}\,\phi_{i+2}(z^{N}_{i})+(-1)^{\,N+1-\varepsilon_{k}}\,\phi_{1}(z^{N}_{N-1})-\delta_{k3}\theta\phi_{2}(\eta_{2})]}{\phi_{1}(z^{N}_{N-1})}\,\phi_{k}(\eta_{k}),\end{aligned}$
with
$\varepsilon_{k}:=\begin{cases}1,&\text{$k$ odd},\\ 0,&\text{$k$ even},\end{cases}\quad\delta_{k3}:=\begin{cases}1,&k=3,\\ 0,&k\neq 3.\end{cases}$
where $A_{i}$ are defined in Appendix B.
  • Proof of Lemma 3.21 :
    From (3.124), we get
    (3.128) $\begin{aligned} \dot{V}_{N}(\eta_{2},z^{N}_{1},\cdots,z^{N}_{N-1})&=\theta\phi_{2}(\eta_{2})\dot{\eta}_{2}+\theta\sum_{i=1}^{N-2}\phi_{i+2}({z^{N}_{i}})\dot{z}^{N}_{i}+\phi_{1}({z^{N}_{N-1}})\dot{z}^{N}_{N-1}.\end{aligned}$
    Our objective is to design a control $u$ such that $\dot{V}_{N}$ satisfies (3.125) thereby guaranteeing dissipation (and hence decay) of the Lyapunov function and the associated closed-loop system. We obtain the expressions for $\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})$ and $\dot{V}_{4}(\eta_{2},z^{4}_{1},\cdots,z^{4}_{3})$ in the form
    (3.129) $\begin{aligned} \dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})=&\phi_{1}(\eta_{1})[\phi_{1}(z^{3}_{2})-\theta\phi_{3}(z^{3}_{1})]-\phi_{1}(z^{3}_{2})\phi_{2}(\eta_{2})+\phi_{3}(\eta_{3})[\theta\phi_{3}(z^{3}_{1})-\theta\phi_{2}(\eta_{2})-\phi_{1}(z^{3}_{2})]+\phi_{1}(z^{3}_{2})(u-u^{*}),\end{aligned}$
    (3.130) $\begin{aligned} \dot{V}_{4}(\eta_{2},z^{4}_{1},\cdots,z^{4}_{3})=&\phi_{1}(\eta_{1})[\phi_{1}(z^{4}_{3})-\theta\phi_{4}(z^{4}_{2})]-\phi_{1}(z^{4}_{3})\phi_{2}(\eta_{2})+\phi_{3}(\eta_{3})[\theta\phi_{3}(z^{4}_{1})-\theta\phi_{2}(\eta_{2})-\theta\phi_{4}(z^{4}_{2})+\phi_{1}(z^{4}_{3})]\\ &+\phi_{4}(\eta_{4})[\theta\phi_{4}(z^{4}_{2})-\theta\phi_{3}(z^{4}_{1})-\phi_{1}(z^{4}_{3})]+\phi_{1}(z^{4}_{3})(u-u^{*}).\end{aligned}$
    For $N=3$ and $N=4$, by introducing the fictitious controls $\eta_{3},\eta_{4},\eta_{1}$ and selecting suitable positive gains $c_{i}$ together with an appropriate control law, we obtain dissipation and guarantee closed-loop stability. Thus, substituting the controls specified in (3.65) and (3.110) into the expressions for $\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})$ and $\dot{V}_{4}(\eta_{2},z^{4}_{1},\cdots,z^{4}_{3})$ and simplifying yields relation (3.125). The constants $c_{i}$ in (3.125) are chosen so that the following conditions are satisfied
    (3.131) $\phi_{3}(z^{N}_{1})=c_{1}\phi_{2}(\eta_{2}),\;\phi_{4}(z^{N}_{2})=c_{2}\phi_{3}(z^{N}_{1}),\;\phi_{5}(z^{N}_{3})=c_{3}\phi_{4}(z^{N}_{2}),\cdots\cdots,\phi_{N}(z^{N}_{N-2})=c_{N-2}\phi_{N-1}(z^{N}_{N-3}),\;\phi_{1}(z^{N}_{N-1})=c_{N-1}\phi_{N}(z^{N}_{N-2}).$
    These constants $c_{i}>0$ are designed to preserve the model’s mechanism and to ensure dissipation of the system along the coordinates $\eta_{2},z^{N}_{1},z^{N}_{2},\dots,z^{N}_{N-1}$. Applying the transformations in (3.123), we obtain the general form of the derivative of the Lyapunov function for any fixed $N\geq 3$